The length of a rectangle is 5 inches more than its width, x. The area of a rectangle can be represented by the equation x2 + 5x = 300. What are the measures of the width and the length?
Width = a0 inches
Length = a1 inches

Answer by stanbon(75887)

The length of a rectangle is 5 inches more than its width, x.
The area of a rectangle can be represented by the equation x2 + 5x = 300. What are the measures of the width and the length?
----
Factor::
x^2 + 5x - 300 = 0
(x-15)(x+20) = 0
----
x = 15 is the width.
x+5 = 20 is the length
-------------
Cheers,
Stan H.
------------